Ticket Number: HEA-NOC/20100420-2 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: NUIG
Ticket Opened: 20100420 09:56 UTC+1 Problem Start: 20100611 22:00 UTC+1
Ticket Update: 20100608 15:31 UTC+1 Problem End: 20100612 08:00 UTC+1
Site/line: CARRAROE, CASTLEBAR, DERI, GTI, IT Athlone, IT Sligo, LYIT,
LYIT-Killybegs, Marine Institute, MARTINRYAN, mhars, NIRAN,
NUIG, NUIG Clinical Sciences, NUIG-Gweedore, PAC, Schools
Network, Sean, St. Angela's
NUIG are expanding their Data Centre. Installation of a new 160KVA UPS
requires a new module to be added to the main Data Centre power
distribution board. This board needs to be powered off in order for the
modification to be made.
UPS and Generator and therefore all power to the data centre, including
HEAnet's PoP, will not be available during the works. The work will take
place over 8 hours, from 09.00-17.00 on Saturday 12th June 2010.
Loss of resiliency
Loss of Connectivity
GMIT (and all sub-sites)
NUIG (and all sub-sites)
-Open Learning Centre
-Sean Scoil Shailerna
-Aras Shorcha Ni Ghuairim
-Aras Mhairtin Ui Chadhain
20100420 09:29 BMA Opening ticket to track work. Clients will be contacted
individually to ensure date of outage is acceptable.
20100511 14:48 CS Since this outage is a problem for IT Sligo and the work
cannot be easily re-scheduled, NUIG will arrange for HEAnet kit to be
powered by the mains from nearby offices during the maintenance work
thereby avoiding a potential 8-hour long outage.
However, as not all of the HEAnet kit has dual power supplies, this will
result in two short outages: the first when the power supply is switched to
the temporary supply before maintenance begins and the second when
switching back after the work is complete.
We are currently waiting to confirm outage times with IT Sligo.
20100520 15:25 BN This work has now been scheduled to commence at 22:00 on
Friday 11th June. A short, ~3m, outage will take place just after 22:00
when replacement power will be put in place. Full power supply will be
swapped back in before 08:00 on the 12th with another short outage.
20100608 15:29 GL Work is to commence as planned on Friday 11th June at
22:00. However, full power supply will now be swapped back on the morning
of the 13th.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20100420-2
HEAnet Limited. Registered in Ireland, No. 275301.